Baker Industries’ net income is $25,000, its interest expense is $4,000, and its tax rate is 40%. Its notes payable equals $27,000, long-term debt equals $70,000, and common equity equals $255,000. The firm finances with only debt and common equity, so it has no preferred stock. What are the firm’s ROE and ROIC? Round your answers to two decimal places. Do not round intermediate calculations.

ROE %
ROIC %
Answer #1
ROE = Net income / Common equity = 25000 / 255000 9.80%
Earnings before tax = Net income / ( 1 - Tax%) = 25000 / ( 1 - 40% ) 41667
Earnings before interest and taxes (EBIT) = Earnings before tax + Interest expense = 41667 + 4000    45667
Total invested capital = Long term debt + Common equity = 70000 + 255000 325000
ROIC = EBIT * ( 1 - Tax% ) / Total invested capital = 45667 * ( 1 - 40% ) / 325000 8.43%